Circular No. 7954

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002el IN NGC 6986 L. Wang, Lawrence Berkeley National Laboratory (LBNL), writes that

preliminary reduction of spectropolarimetry (range 333–860 nm; spectral resolution 1.27 nm) of SN 2002el (cf. IAUC 7953), obtained with the European Southern Observatory’s Very Large Telescope UT3 (+ FORS1) on Oct. 14, shows it to be a type-Ia supernova around optical maximum with < 0.2 percent of intrinsic polarization.

F. Patat, European Southern Observatory; and P. Meikle, Imperial College, London, on behalf of the European Research and Training Network on the Physics of Type-Ia Supernovae, report that spectrograms (ranges 390–500 and 600–710 nm; resolution 0.17 nm) of SN 2002el were obtained at Calar Alto Observatory by J. Aceituno on Aug. 14.97 UT at the 3.5m telescope (+ TWIN double-channel spectrograph). A full reduction of the two 1800-s spectra shows the object to be a type-Ia supernova. A comparison with the spectra of SN 1994D (Patat et al. 1996, MNRAS 278,

111) indicates that SN 2002el was observed 8–10 days before maximum light. The recession velocity of the parent galaxy, deduced from the match to the spectra of SN 1994D, is 9000 ± 1000 km/s.

W. M. Wood-Vasey, University of California at Berkeley (UCB) and LBNL, reports that prediscovery NEAT observations of SN 2002el, taken with the Oschin 1.2-m telescope at Palomar Observatory on Aug. 8.29, 8.31, and 8.36 UT, clearly show the new object at unfiltered mag 19.5. Images taken with the same telescope on July 17.30 and 17.87 show nothing at this position (limiting mag 20.5).

SUPERNOVA 2002dz IN MCG –01-1-52 A. V. Filippenko, R. J. Foley, M. Papenkova, and B. Swift, UCB, re-

port that inspection of CCD spectra (range 310–1000 nm), obtained on Aug. 10 UT with the Shane 3-m reflector at Lick Observatory, reveals that SN 2002dz (IAUC 7935) is of type Ib or Ic, well into the nebular phase. Strong, broad emission lines of [O I], [Ca II], and Ca II (near-infrared triplet) are present, with no evidence for broad Hα. The [O I] and [Ca II] line widths (FWHM ∼ 5400 km/s) are smaller than average for nebular type-Ic supernovae (Matheson et al. 2001, A.J. 121, 1648), perhaps suggesting that the object is actually of type Ib, with more massive ejecta than are present with typical type-Ic supernovae.

2002 August 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green